OpenStudy (anonymous):

Determine the vertex by completing the square. f(x)2x^2+12x+16.

OpenStudy (anonymous):

start with \[f(x)=2(x^2+6x)+16\] then \[f(x)=2(x+3)^2+k\] do find \(k\) find \(f(-3)\)

OpenStudy (anonymous):

*to find \(k\)